Please help me solve this question: 4^x=17
Answer by stanbon(75887)

4^x=17
-----
Whenever you have the variable in the exponent, take the log or ln.
-------
log4^x = log17
x*log(4) = log(17)
x = log(17)/log(4)
x = 2.0437
================
Cheers,
Stan H.
================